A brass rod is 60.04cm at 20ºC. To what temperature must it be raised to make it 60.05cm long?

1
Expert's answer
2022-05-02T13:42:13-0400
"l=l_0(1+\alpha\Delta T)"

"\Delta T=\frac{1}{\alpha}(l/l_0-1)"

"\Delta T=\frac{1}{19*10^{-6}}(60.05/60.04-1)=8.8^\circ\rm C"

"T=20^\circ\rm C+8.8^\circ\rm C=28.8^\circ\rm C"